7. (4 pts) Graph the following linear equation on the grid below: y = -x-5 pe is 3 3
Answer #1

Given that y=2x-5 at X=-3 5 Y = {(-3)–5=-9-10 .-9.5 2 x = -2 ; G=(-2)-5--3-52-8 3/2 (-1)-5 x = -1 y = = -3-10 =-6.5 22 = 0 y x=1 y = 3(0)-5--5 y = {(1)-5= 3,710 = -3.5 y = {(2)-5= -2 y = 3 (3)-5= 9-10 x = 2 De = 3 =-0.5 2상 -3. 0 - 2-1 컷 1 2 ) - 2 따지 - 3 - 9. Hors) - - 2 - 8 여 (-3, -9.s)